when a truck is moving at a speed of 10ms. when the driver seed charge in front, he slow and decelerated at -1.5ms^_2. Him reaction time to hit the break was 0.5sec. will the truck hit the child?


1
Expert's answer
2021-01-21T09:57:56-0500

Initial velocity, "u = 10 m\/s"

Reaction time, "t = 0.5 s"

Retardation, "a = 1.5 m\/s^2"

Let x be the distance of the boy from the truck.


To avoid accident,

"ut +(uT-\frac{1}{2}aT^2) < x"


"(10 \times 0.5) + (10T-\frac{1}{2}(1.5)T^2) < x"


where T is the time after reaction time to time to reach to object.